patrick drove 6 hours to attend a meeting. if his return speed was 12 mph less, and took 8 hours, what were his two speeds
I'm in pre-calc and cant remember how to help my brother with this algebra problem help would be great:)

Answer by lwsshak3(11628)

patrick drove 6 hours to attend a meeting. if his return speed was 12 mph less, and took 8 hours, what were his two speeds
***
let x=speed to meeting
x-12=return speed
Distance=speed*travel time
...
6x=8(x-12)
6x=8x-96
2x=96
x=48
x-12=36
speed to meeting=48 mph
return speed=36 mph